# A DRL-Empowered Multi-Level Jamming Approach for Secure Semantic Communication

### v1

@@ I Introduction
Semantic communication (SemCom) has emerged as a promising paradigm that improves communication efficiency by transmitting only task-relevant information, rather than raw symbols. Significant progress has been made in various data modalities, including text [1], image [2], speech [3], and video [4]. However, by focusing primarily on the essential semantics, SemCom typically eliminates traditional channel coding, which adds redundancy to improve reliability and provide implicit security. The absence of this redundancy introduces a vulnerability: transmitted data becomes more susceptible to eavesdropping, as intercepted signals can directly reveal sensitive content. As a result, secure SemCom over wiretap channels has become a crucial research direction and has recently attracted considerable attention [5]. Existing approaches can be broadly categorized into three categories: adversarial training-based methods, encryption-based methods, and physical layer-based methods.
From the adversarial training perspective, Erdemir et al. [6] proposed a VAE-based joint source-channel coding (JSCC) scheme that enhances security by maximizing the mutual information (MI) between the source and the legitimate channel output while penalizing information leakage. However, its security relies on variational approximations of MI, which may reduce its robustness against stronger adversaries. From the key-sharing-based encryption perspective, Luo et al. [7] developed an encrypted SemCom system (ESCS) that integrates a semantics-based symmetric cryptosystem learned through neural networks (NNs), and further proposed an adversarial encryption training strategy. However, it requires a secure channel for key exchange, which remains a limitation for practical deployment. More recent efforts have explored the use of homomorphic encryption in SemCom. Meng et al. [8] showed that semantic features can be retained in homomorphically encrypted data and proposed a privacy-preserved DeepJSCC framework, where activation and pooling functions are modified to support homomorphic operations.
@@ II System Model
In addition to the source image $\textbf{X}\in\mathbb{R}^{H\times W\times C}$, other data is also included as input to the transmitter. We denote the entire input as Z. Alice employs a transmitter to encode the input Z in order to protect X, generating the semantic information to be transmitted, denoted as $\textbf{Y}\in\mathbb{R}^{N_{m}\times L_{c}}$:
@@
Subsequently, Y is normalized to satisfy the average transmit power constraint $P$, resulting in the normalized signal $\tilde{\textbf{Y}}$. This signal $\tilde{\textbf{Y}}$ is then transmitted through MIMO Rayleigh fading channels to Bob. Meanwhile, Eve intercepts the transmitted semantic information through her own channels. In this paper, we assume that the channel conditions for both the legitimate user and the eavesdropper are comparable. Accordingly, the semantic information received by Bob and Eve are denoted as ${\textbf{Y}}_{1}\in\mathbb{R}^{N_{n}\times L_{c}}$ and ${\textbf{Y}}_{2}\in\mathbb{R}^{N_{n}\times L_{c}}$, respectively, and are expressed as
@@
where $\textbf{H}\in\mathbb{R}^{N_{n}\times N_{m}}$ denotes the channel matrix between Alice and Bob/Eve. Considering a conventional setting, we assume $N_{m}=N_{n}$. The channel coefficients h follow the distribution $\textbf{h}\sim\mathcal{N}(0,\textbf{I}_{N_{n}})$, where $\textbf{I}_{N_{n}}$ denotes the $N_{n}\times N_{n}$ identity matrix. $\textbf{N}_{1/2}\in\mathbb{R}^{N_{n}\times L_{c}}$ represents the additive white Gaussian noise (AWGN) term for Bob/Eve, where each element follows a Gaussian distribution $\mathcal{N}(0,\sigma_{1/2}^{2})$. The channel SNR between Alice and Bob/Eve can be calculated by $\mathrm{SNR}_{\rm leg/\rm eve}=10\log_{10}\left(\frac{P}{\sigma_{1/2}^{2}}\right)\,(\mathrm{dB})$.
At the receiver, before feeding the received semantic information ${\textbf{Y}}_{1}$ and ${\textbf{Y}}_{2}$ into their respective semantic decoders, Bob and Eve first perform signal detection on the received signals using minimum mean-squared error (MMSE) equalization [14]. MMSE equalization operates based on the estimated CSI $\hat{\textbf{H}}\in\mathbb{R}^{N_{n}\times N_{m}}$. In this paper, the estimated CSI is assumed to be perfect. The MMSE equalization process is as follows:
$\hat{\textbf{Y}}_{1/2}={\hat{\textbf{H}}}^{H}\left(\hat{\textbf{H}}{\hat{\textbf{H}}}^{H}+\frac{\sigma_{1/2}^{2}}{P}\textbf{I}\right)^{-1}\textbf{Y}_{1/2}.$ (3)
Here, $\hat{\textbf{Y}}_{1}$ and $\hat{\textbf{Y}}_{2}$ denote the MMSE-equalized versions of ${\textbf{Y}}_{1}$ and ${\textbf{Y}}_{2}$, respectively. $\sigma_{1/2}^{2}$ denotes the variance of the AWGN, and $P$ denotes the transmit power constraint.
@@ III-A Transmitter
After generating the semantic information and multi-level jamming signals, Alice applies the precoding matrices $\textbf{V}_{1}$, $\textbf{V}_{2}$, and $\textbf{V}_{3}$, which are produced by the DDPG agent (to be discussed later), to precode the signals $\textbf{S}_{1}$, $\textbf{S}_{2}$, and $\textbf{S}_{3}$, respectively. These precoded signals are then combined through superposition coding to form the transmitted semantic information, denoted as Y. The process is described by the following equation:
@@ III-C The DDPG Agent
    $\begin{split}\textbf{s}_{t}=\left[{\rm SNR}^{\rm leg}_{t},{\rm SNR}^{\rm eve}_{t},CU,\right.\\ \left.{\rm PSNR}^{\rm leg}_{t-1},{\rm PSNR}^{\rm eve}_{t-1},\mathcal{L}_{t-1},t/T\right],\end{split}$ (12)
    where ${\rm SNR}^{\rm leg}_{t}$ and ${\rm SNR}^{\rm eve}_{t}$ denote the SNRs of the legitimate user and the eavesdropper, respectively. $CU$ controls the compression ratio of the system. ${\rm PSNR}^{\rm leg}_{t-1}$ and ${\rm PSNR}^{\rm eve}_{t-1}$ indicate the image reconstruction quality (PSNR) from the previous time step for Bob and Eve, respectively. $\mathcal{L}_{t-1}$ denotes the previous loss, and $t/T$ represents the relative training progress.
@@ III-D1 Stage 1
In this training stage, we focus on training the semantic encoder $f_{\text{SE}}$ and the semantic decoder of the legitimate user $f_{\text{SD1}}$, while keeping all other modules frozen. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{S}_{1}$. The system is optimized using
@@ III-D2 Stage 2
In this training stage, multi-level jamming is introduced into the system, but the DDPG agent is not yet involved. All modules are activated except for the eavesdropper’s decoder and the DDPG agent. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{S}_{1}+\textbf{S}_{2}+\textbf{S}_{3}$. The loss function remains:
@@ III-D3 Stage 3
In this training stage, the DDPG agent remains inactive. The eavesdropper’s decoder $f_{\text{SD2}}$ is trained independently, with all other modules kept frozen. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{S}_{1}+\textbf{S}_{2}+\textbf{S}_{3}$. The loss function used is:
@@ III-D4 Stage 4
In this training stage, the DDPG agent is introduced and jointly trained with the SemCom model (environment) through an alternating update scheme. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{V}_{1}\textbf{S}_{1}+\textbf{V}_{2}\textbf{S}_{2}+\textbf{V}_{3}\textbf{S}_{3}$. Specifically, the training alternates between two stages:
  • Agent update stage: At the beginning of every $K$ epochs, the DDPG agent observes the current environment state and generates a new action $\textbf{a}_{t}$, corresponding to a set of precoding matrices $\{\textbf{V}_{1},\textbf{V}_{2},\textbf{V}_{3}\}$. This action remains fixed throughout the subsequent $K$ epochs of environment training. After receiving the reward and state transition at the end of the $K$ epochs, the agent stores the resulting experience tuple in its replay buffer. If the buffer contains enough samples, the agent updates its actor and critic networks via mini-batch gradient descent.
@@ III-D5 Stage 5
In this training stage, the optimal policy obtained in stage 4, selected based on the highest reward, is saved and directly applied to guide both the training and testing of the final system. The transmitted semantic information is generated in the same way as in stage 4, and the same loss function is used.
@@ IV-A2 Training Settings
We set $N_{m}=N_{n}=4$. Both users are assumed to experience the same channel SNR, given by $\mathrm{SNR}_{\mathrm{leg}}=\mathrm{SNR}_{\mathrm{eve}}\in\{0,5,10,15,20\}\,\mathrm{dB}$. The value of compression ratio (CR) is controlled by the variable $CU$. When $CU=1$, the corresponding CR is $1/96$. The CR takes values from the set $CR\in\{1/96,2/96,3/96,4/96,5/96\}$.
@@
The learning rate in the first three training stages is set to $1\times 10^{-3}$ and maintained for 100 epochs. In the fourth training stage, the learning rate is set to $5\times 10^{-4}$ and maintained for $T\cdot K$ epochs, which correspond to 7500 epochs. In the fifth training stage, the learning rate is set to $2\times 10^{-4}$ and maintained for 200 epochs.
@@ IV-A3 The Benchmarks
The first benchmark is the ESCS scheme [7], which is a key-based encryption approach designed for secure SemCom systems over wiretap channels. The second benchmark is the EJ scheme [13], which is a jamming-based scheme developed for secure communication systems over MIMO wiretap channels. Both benchmarks employ the classical singular value decomposition (SVD) precoding and MMSE equalization.
@@ IV-B1 Different Channel SNRs
Figure 2: Comparison of PSNR performance for both users under different channel SNRs using different approaches. The compression ratio is fixed at 1/96, with $N_{m}$ and $N_{n}$ both set to 4. We set $\rm SNR_{leg}=\rm SNR_{eve}$.
@@
Figure 3: Comparison of PSNR performance for both users under different compression ratios using different approaches. We set $\rm SNR_{leg}=\rm SNR_{eve}=10dB$, with $N_{m}$ and $N_{n}$ both set to 4.
@@ IV-B2 Different Compression Ratios
In this subsection, we evaluate the PSNR performance of the proposed system under different compression ratios and compare it with the two benchmarks. As shown in Fig. 3, the PSNR of the legitimate user increases steadily as the compression ratio grows. Among all evaluated methods, the proposed method consistently achieves the highest reconstruction quality across all compression ratios. Specifically, at a compression ratio of 5/96, our method attains a PSNR of 20.92 dB, outperforming EJ (20.14 dB) and ESCS (20.54 dB). In contrast to the legitimate user, the eavesdropper’s PSNR exhibits a decreasing trend with higher compression ratios. At the highest compression ratio, our method reduces the eavesdropper’s PSNR to 6.03 dB, which is lower than that of EJ (6.46 dB) and ESCS (6.20 dB), indicating improved system security. Overall, compared with the benchmarks, the proposed system consistently achieves a superior trade-off between task performance and system security, demonstrating high effectiveness and robustness under varying compression ratios.
@@ V Conclusion
In this paper, we proposed a SemCom framework that integrated a DRL-empowered multi-level jamming strategy to enhance transmission security over MIMO fading wiretap channels. The framework was designed for a realistic and challenging scenario where both the legitimate user and the eavesdropper experienced comparable channel conditions, without relying on secret key exchange or any prior shared knowledge. Specifically, two levels of jamming signals were introduced: semantic layer jamming, generated by encoding task-irrelevant text data, and physical layer jamming, generated by encoding random Gaussian noise. These jamming signals were superposed with task-relevant semantic information to effectively obscure the transmitted semantics from the eavesdropper while maintaining high reconstruction quality for the legitimate user. To optimize and empower this process, we employed the DDPG algorithm to design and adapt the precoding matrices for both task-relevant semantic information and multi-level jamming signals. This approach effectively balanced the trade-off between the legitimate user’s reconstruction quality and overall system security. Moreover, the DDPG agent operated only during training, introducing no additional computational overhead during inference. Experimental results demonstrated that, compared with the ESCS and EJ benchmarks, our approach achieved comparable security while improving the legitimate user’s PSNR by up to approximately 0.6 dB. These results confirmed that the proposed DRL-empowered multi-level jamming framework provided an effective and practical solution for secure SemCom systems operating under challenging conditions.

### v3

@@ I Introduction
Semantic communication (SemCom) has emerged as a promising paradigm that improves communication efficiency by transmitting only task-relevant information, rather than raw symbols. Significant progress has been made in various data modalities, including text [11], image [2], speech [5], and video [7]. However, by focusing primarily on the essential semantics, SemCom typically eliminates traditional channel coding, which adds redundancy to improve reliability and provide implicit security. The absence of this redundancy introduces a vulnerability: transmitted data becomes more susceptible to eavesdropping, as intercepted signals can directly reveal sensitive content. As a result, secure SemCom over wiretap channels has become a crucial research direction and has recently attracted considerable attention [4]. Existing approaches can be broadly categorized into three groups: adversarial training-based methods, encryption-based methods, and physical layer-based methods.
From the adversarial training perspective, Erdemir et al. [6] proposed a variational autoencoder (VAE)-based deep joint source-channel coding (DeepJSCC) scheme that enhances security by maximizing the mutual information (MI) between the source and the legitimate channel output while penalizing information leakage. However, its security relies on variational approximations of MI, which may reduce its robustness against stronger adversaries. From the key-sharing-based encryption perspective, Luo et al. [10] developed an encrypted SemCom system (ESCS) that integrates a semantics-based symmetric cryptosystem learned through neural networks (NNs), and further proposed an adversarial encryption training strategy. However, it requires a secure channel for key exchange, which remains a limitation for practical deployment. More recent efforts have explored the use of homomorphic encryption in SemCom. Meng et al. [12] showed that semantic features can be retained in homomorphically encrypted data and proposed a privacy-preserved DeepJSCC framework, where activation and pooling functions are modified to support homomorphic operations.
@@ II System Model
In addition to the source image $\textbf{X}\in\mathbb{R}^{H\times W\times C}$, other information is also included as input to the transmitter. We denote the entire input as Z. Alice employs a transmitter to encode the input Z in order to protect X, generating the semantic information to be transmitted, denoted as $\textbf{Y}\in\mathbb{R}^{N_{m}\times L_{c}}$:
@@
Subsequently, Y is normalized with respect to its average power in order to satisfy the transmit power constraint $P$ and then converted into a complex-valued representation, resulting in the normalized signal $\tilde{\textbf{Y}}\in\mathbb{C}^{N_{m}\times\frac{L_{c}}{2}}$. This signal $\tilde{\textbf{Y}}$ is then transmitted through MIMO Rayleigh fading channels to Bob. Meanwhile, Eve intercepts the transmitted semantic information through her own channels. Accordingly, the semantic information received by Bob and Eve are denoted as ${\textbf{Y}}_{1}\in\mathbb{C}^{N_{n}\times\frac{L_{c}}{2}}$ and ${\textbf{Y}}_{2}\in\mathbb{C}^{N_{n}\times\frac{L_{c}}{2}}$, respectively, and are expressed as
@@
where $\textbf{H}\in\mathbb{C}^{N_{n}\times N_{m}}$ denotes the channel matrix between Alice and Bob/Eve. Considering a conventional setting, we assume $N_{m}=N_{n}$. The channel coefficients h follow the distribution $\textbf{h}\sim\mathcal{CN}(0,\textbf{I}_{N_{n}})$, where $\textbf{I}_{N_{n}}$ denotes the $N_{n}\times N_{n}$ identity matrix. $\textbf{N}_{1/2}\in\mathbb{C}^{N_{n}\times\frac{L_{c}}{2}}$ represents the additive white Gaussian noise (AWGN) term for Bob/Eve, where each element follows a complex Gaussian distribution $\mathcal{CN}(0,\sigma_{1/2}^{2})$. The channel SNR between Alice and Bob/Eve can be calculated by $\mathrm{SNR}_{\rm leg/\rm eve}=10\log_{10}\left(\frac{P}{\sigma_{1/2}^{2}}\right)\,(\mathrm{dB})$.
At the receiver, before feeding the received semantic information ${\textbf{Y}}_{1}$ and ${\textbf{Y}}_{2}$ into their respective semantic decoders, Bob and Eve first perform signal detection on the received signals using minimum mean-squared error (MMSE) equalization [15]. MMSE equalization operates based on the estimated channel state information (CSI) $\hat{\textbf{H}}\in\mathbb{C}^{N_{n}\times N_{m}}$. In this paper, the estimated CSI is assumed to be perfect. The MMSE equalization process is as follows:
$\bar{\textbf{Y}}_{1/2}={\hat{\textbf{H}}}^{H}\left(\hat{\textbf{H}}{\hat{\textbf{H}}}^{H}+\frac{\sigma_{1/2}^{2}}{P}\textbf{I}\right)^{-1}\textbf{Y}_{1/2}.$ (3)
Here, $\bar{\textbf{Y}}_{1}$ and $\bar{\textbf{Y}}_{2}$ denote the MMSE-equalized versions of ${\textbf{Y}}_{1}$ and ${\textbf{Y}}_{2}$, respectively. $\sigma_{1/2}^{2}$ denotes the variance of the AWGN, and $P$ denotes the transmit power constraint. Subsequently, $\bar{\textbf{Y}}_{1}$ and $\bar{\textbf{Y}}_{2}$ are converted back to real-valued representations, denoted by $\hat{\textbf{Y}}_{1}$ and $\hat{\textbf{Y}}_{2}$.
@@ III-A Transmitter
After generating the semantic information and multi-level jamming signals, Alice applies the precoding matrices $\textbf{V}_{1}$, $\textbf{V}_{2}$, and $\textbf{V}_{3}$, which are produced by the deep deterministic policy gradient (DDPG) agent (to be discussed later), to precode the signals $\textbf{S}_{1}$, $\textbf{S}_{2}$, and $\textbf{S}_{3}$, respectively. These precoded signals are then combined through superposition coding to form the transmitted semantic information, denoted as Y. The process is described by the following equation:
@@ III-C The DDPG Agent
    $\begin{split}\textbf{s}_{t}=\left[{\rm SNR}^{\rm leg}_{t},{\rm SNR}^{\rm eve}_{t},CU,\right.\\ \left.{\rm PSNR}^{\rm leg}_{t-1},{\rm PSNR}^{\rm eve}_{t-1},\mathcal{L}_{t-1},\frac{t}{T}\right],\end{split}$ (12)
    where ${\rm SNR}^{\rm leg}_{t}$ and ${\rm SNR}^{\rm eve}_{t}$ denote the SNRs of the legitimate user and the eavesdropper, respectively. $CU$ controls the compression ratio (CR) of the system. ${\rm PSNR}^{\rm leg}_{t-1}$ and ${\rm PSNR}^{\rm eve}_{t-1}$ indicate the image reconstruction quality (PSNR) from the previous time step for Bob and Eve, respectively. $\mathcal{L}_{t-1}$ denotes the previous loss, and $\frac{t}{T}$ represents the relative training progress.
@@ III-D1 Stage 1
In this stage, we focus on training the semantic encoder $f_{\text{SE}}$ and the semantic decoder of the legitimate user $f_{\text{SD1}}$, while keeping all other modules frozen. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{S}_{1}$. The system is optimized using
@@ III-D2 Stage 2
In this stage, multi-level jamming is introduced into the system, but the DDPG agent is not yet involved. All modules are activated except for the eavesdropper’s decoder and the DDPG agent. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{S}_{1}+\textbf{S}_{2}+\textbf{S}_{3}$. The loss function remains:
@@ III-D3 Stage 3
In this stage, the DDPG agent remains inactive. The eavesdropper’s decoder $f_{\text{SD2}}$ is trained independently, with all other modules kept frozen. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{S}_{1}+\textbf{S}_{2}+\textbf{S}_{3}$. The loss function used is:
@@ III-D4 Stage 4
In this stage, the DDPG agent is introduced and jointly trained with the SemCom model (environment) through an alternating update scheme. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{V}_{1}\textbf{S}_{1}+\textbf{V}_{2}\textbf{S}_{2}+\textbf{V}_{3}\textbf{S}_{3}$. Specifically, the training alternates between two stages:
  • Agent update stage: At the beginning of each $K$-epoch interval, the DDPG agent observes the current environment state and generates a new action $\textbf{a}_{t}$, corresponding to a set of precoding matrices $\{\textbf{V}_{1},\textbf{V}_{2},\textbf{V}_{3}\}$. This action remains fixed throughout the subsequent $K$ epochs of environment training. After receiving the reward and state transition at the end of the $K$ epochs, the agent stores the resulting experience tuple in its replay buffer. If the buffer contains enough samples, the agent updates its actor and critic networks via mini-batch gradient descent.
@@ III-D5 Stage 5
In this stage, the optimal policy obtained in stage 4, selected based on the highest reward, is saved and directly applied to guide both the training and testing of the final system. The transmitted semantic information is generated in the same way as in stage 4, and the same loss function is used.
@@ IV-A2 Training Settings
We set $N_{m}=N_{n}=4$. Both users are assumed to experience the same channel SNR, given by $\mathrm{SNR}_{\mathrm{leg}}=\mathrm{SNR}_{\mathrm{eve}}\in\{0,5,10,15,20\}\,\mathrm{dB}$. The CR is controlled by the variable $CU$. When $CU=1$, the corresponding CR is $1/96$. The CR takes values from the set $CR\in\{1/96,2/96,3/96,4/96,5/96\}$.
@@
The learning rate in the first three training stages is set to $1\times 10^{-3}$ and maintained for 100 epochs. In the fourth training stage, the learning rate is set to $5\times 10^{-4}$ and maintained for $T\cdot K$ epochs, which corresponds to 7500 epochs. In the fifth training stage, the learning rate is set to $2\times 10^{-4}$ and maintained for 200 epochs.
@@ IV-A3 The Benchmarks
The first benchmark is the ESCS scheme [10], which is a key-based encryption approach designed for secure SemCom systems over wiretap channels. The second benchmark is the EJ scheme [14], which is a jamming-based scheme developed for secure communication systems over MIMO wiretap channels. In both benchmarks, we employ classical singular value decomposition (SVD) precoding and MMSE equalization.
@@ IV-B1 Different Channel SNRs
Figure 2: Comparison of PSNR performance for both users under different channel SNRs using different approaches. The CR is fixed at $1/96$, with $N_{m}$ and $N_{n}$ both set to 4. We set $\rm SNR_{leg}=\rm SNR_{eve}$.
@@
Figure 3: Comparison of PSNR performance for both users under different CRs using different approaches. We set $\rm SNR_{leg}=\rm SNR_{eve}=10dB$, with $N_{m}$ and $N_{n}$ both set to 4.
@@ IV-B2 Different Compression Ratios
In this subsection, we evaluate the PSNR performance of the proposed system under different CRs and compare it with the two benchmarks. As shown in Fig. 3, the PSNR of the legitimate user increases steadily as the CR grows. Among all evaluated methods, the proposed method consistently achieves the highest reconstruction quality across all CRs. Specifically, at a CR of $5/96$, our method attains a PSNR of 20.92 dB, outperforming EJ (20.14 dB) and ESCS (20.54 dB). In contrast to the legitimate user, the eavesdropper’s PSNR exhibits a decreasing trend with higher CRs. At the highest CR, our method reduces the eavesdropper’s PSNR to 6.03 dB, which is lower than that of EJ (6.46 dB) and ESCS (6.20 dB), indicating improved system security. Overall, compared with the benchmarks, the proposed system consistently achieves a superior trade-off between task performance and system security, demonstrating high effectiveness and robustness under varying CRs.
@@ V Conclusion
In this paper, we proposed a SemCom framework that integrated a DRL-empowered multi-level jamming strategy to enhance transmission security over MIMO fading wiretap channels. The framework was designed for a realistic and challenging scenario where both the legitimate user and the eavesdropper experienced comparable channel conditions, without relying on secret key exchange or any prior shared knowledge. Specifically, two levels of jamming signals were introduced: semantic layer jamming, generated by encoding task-irrelevant text data, and physical layer jamming, generated by encoding random Gaussian noise. These jamming signals were superposed with task-relevant semantic information to effectively obscure the transmitted semantics from the eavesdropper while maintaining high reconstruction quality for the legitimate user. To optimize and empower this process, we employed the DDPG algorithm to design and adapt the precoding matrices for both task-relevant semantic information and multi-level jamming signals. This approach effectively balanced the trade-off between the legitimate user’s reconstruction quality and overall system security. Moreover, the DDPG agent operated only during training, introducing no additional computational overhead during inference. Experimental results demonstrated that, compared with the ESCS and EJ benchmarks, our approach achieved comparable security while improving the legitimate user’s PSNR by up to approximately 0.6 dB. These results confirm that the proposed DRL-empowered multi-level jamming framework provides an effective and practical solution for secure SemCom systems operating under challenging conditions.